### v1

# Entangled in Representations: Mechanistic Investigation of Cultural Biases in Large Language Models

Abstract
The growing deployment of large language models (LLMs) across diverse cultural contexts necessitates a better understanding of how the overgeneralization of less documented cultures within LLMs’ representations impacts their cultural understanding. Prior work only performs extrinsic evaluation of LLMs’ cultural competence, without accounting for how LLMs’ internal mechanisms lead to cultural (mis)representation. To bridge this gap, we propose Culturescope, the first mechanistic interpretability-based method that probes the internal representations of LLMs to elicit the underlying cultural knowledge space. CultureScope utilizes a patching method to extract the cultural knowledge. We introduce a cultural flattening score as a measure of the intrinsic cultural biases. Additionally, we study how LLMs internalize Western-dominance bias and cultural flattening, which allows us to trace how cultural biases emerge within LLMs. Our experimental results reveal that LLMs encode Western-dominance bias and cultural flattening in their cultural knowledge space. We find that low-resource cultures are less susceptible to cultural biases, likely due to their limited training resources. Our work provides a foundation for future research on mitigating cultural biases and enhancing LLMs’ cultural understanding. Our codes and data used for experiments are publicly available [^1] https://github.com/copenlu/CultureScope.
1 Introduction
Large language models (LLMs) are increasingly being used in culturally diverse contexts, where understanding and responding appropriately to each culture is essential (Pandya and Holia, 2023; Salemi et al., 2023; Liu et al., 2025). However, the cultural knowledge acquired by LLMs is largely shaped by the data they are trained on, which is predominantly Western-centric (Santurkar et al., 2023). This results in severe geographic imbalances, where some regions receive disproportionate attention while others remain marginalized (Nguyen et al., 2023a). While LLMs do acquire some cross-cultural patterns during training (Hershcovich et al., 2022; Arora et al., 2023), this learning often results in overgeneralization. For example, when asked about a popular leisure activity for retired men in Azerbaijan, GPT-4 responded with chess—a plausible answer, but one that reflects general stereotypes of post-Soviet or Eastern European regions rather than Azerbaijan specifically (Myung et al., 2024).
Figure 1: Given the question about the popular domestic vacation spot in Greece, Culturescope first generates an answer to cultural Question at the Inference stage. Then, it reads the hidden representation from the Inference stage and elicits the cultural knowledge used for ‘Zakynthos’ at the Scoping-in stage. We finalize a list of cultural knowledge after the Filtering stage. Culturescope unveils the internal mechanism of LLMs that cannot be revealed through the Inference stage alone.
However, these patterns of overgeneralization directly perpetuate cultural biases such as cultural flattening. Measuring such overgeneralizations through model outputs alone proves insufficient since generated outputs mask the underlying mechanisms that lead to cultural (mis)representation. To facilitate the examination of the underlying mechanisms, we propose to study LLMs’ cultural understanding with mechanistic interpretability (MI) techniques. MI techniques provide us with methods that can directly examine how cultural biases discovered by the extrinsic evaluation Santurkar et al. (2023) is internally processed within model representations, revealing where and how harmful generalizations emerge. We are the first to propose an approach for intrinsic cultural bias evaluation.
In this work, we introduce Culturescope, a method to probe internal representations and surface the cultural knowledge activated during cultural understanding tasks (§4.1). By surfacing this knowledge, we reveal the internal cultural knowledge space from which model outputs are generated. Figure 1 illustrates an overview of Culturescope. To examine the intrinsic ‘cultural flattening’ embedded in the parameter space, we introduce a cultural flattening (CF) score, which quantifies the degree of intersection between cultural knowledge decoded by Culturescope (§4.2). We implement our framework on two cultural understanding tasks, cultural commonsense Question Answering (QA) and extractive QA, across three different LLMs.
We further challenge the model’s cultural understanding by creating multiple-choice questions (MCQs) with hard negatives (§3.2.3). Culturally nuanced answers from high-resource cultures or geographically proximate countries are selected as hard negative options to simulate the Western-dominance bias and the cultural flattening. This setup refrains LLMs to leverage surface-level elimination strategies based on the overgeneralization Khan et al. (2025). Analyzing selected options by LLMs with the attention map method Yuksekgonul et al. (2024) allows us to examine whether extrinsic and intrinsic cultural biases align, by revealing which options the model internally attends to (§5.1).
High CF scores of Western and high-resource cultures show that LLMs encode Western-dominance bias and cultural flattening in their cultural knowledge space (§6.2). Model performances on MCQs with hard negatives demonstrate that low-resource cultures are less likely to be affected by the biases, likely due to their limited training resource (§6.3). This implies that LLMs struggle with low-resource cultures due to the lack of parametric knowledge. In §6.4, we find that LLMs over-attend to tokens from Western and high-resource cultures. This suggests Western-dominance bias is more internalized than the cultural flattening. Our work provides useful signal to future research on mitigating the internalized patterns for biases to build a better culturally aligned model.
2 Related Work
Evaluating Cultural Understanding of LLMs
Previous work has proposed evaluation datasets and frameworks to assess LLMs’ cultural understanding ability acquired during pre-training Keleg and Magdy (2023); Naous and Xu (2025); Pawar et al. (2025). BLEnD Myung et al. (2024) provides a multilingual commonsense QA dataset spanning 16 countries and regions, designed to uncover cross-cultural disparities in everyday knowledge. CAMeL Naous et al. (2024) compares LLM behavior in Arabic versus Western settings across tasks like story generation, NER, and sentiment analysis, exposing systematic cultural biases in LLMs. Other multilingual benchmarks Zhou et al. (2025); Hasan et al. (2025); Wang et al. (2024); Cao et al. (2024) construct culturally localized evaluation datasets that span domains such as cuisine, proverbs, news, and reasoning. Across these datasets, performance gaps are consistently observed between high-resource and underrepresented languages and cultures, often linked to pre-training data imbalances that favor dominant regions Naous and Xu (2025).
While these efforts highlight important cross-cultural disparities, they perform an extrinsic evaluation, overlooking the underlying mechanism and cultural knowledge space embedded in LLMs. To address this gap, our paper aims to reveal how culture is embedded, entangled, or flattened within the models’ inner representations.
Mechanistic Interpretability
MI techniques are developed to explain the inner workings of LLMs by identifying responsible model components, such as neurons and attention heads Meng et al. (2022); Geva et al. (2023); Yu et al. (2024). Leveraging their transparency, recent studies have employed MI techniques to investigate how specific behaviors emerge in LLMs. For instance, they have been used to uncover and manipulate components associated with social biases, enabling both diagnostic and steering interventions Liu et al. (2024); Durmus et al. (2024); Yang et al. (2024). Despite growing interest in the cultural capabilities of LLMs, no prior work has explored cultural understanding through the lens of MI. Our study fills this gap by applying MI techniques to probe the internal representation of cultural knowledge in LLMs, offering new insights into how cultural understanding is encoded and organized within the model.
3 Experimental Setup
3.1 Preliminaries
A dataset $D=[(q_{1},C_{1},a_{1}),...,(q_{N},C_{N},a_{N})]$ consists of $N$ tuple instances containing: a question $q$, an option list $C=[c_{0},c_{1},c_{2},a]$ containing three options for MCQ, one gold answer $a$ and a country of interest $y$. For MCQ, an LLM is given $C$ and $q$ to generate an output $O=[o_{0},...,o_{P}]$ consisting of $P$ tokens. For open-ended QA, an LLM is only given $q$ to generate an output $O$.
To generate an answer, an LLM converts a tokenized input text $T=[t_{0},...,t_{S}]$ containing $S$ number of tokens into $d$-dimenstional vectors using the embedding matrix $E\in\mathbb{R}^{|\mathbb{V}\times d|}$. Then, the vectors are processed through $L$ layers, each containing a multi-head self-attention (MHSA) layer and an MLP layer. The hidden representation $x_{i}^{l}$ from a layer $l$, on a token $t_{i}$ is computed by:
$x_{i}^{l}=x_{i}^{l-1}+a_{i}^{l}+m_{i}^{l}$ (1)
where $a_{i}^{l}$ is an output from the MHSA layer and $m_{i}^{l}$ from the MLP layer. The hidden representation from the last layer $x_{i}^{L}$ is converted into a token by calculating the logits with the unembedding layer.
3.2 Datasets
3.2.1 Cultural QA Datasets
We select BLEnD Myung et al. (2024), a cultural commonsense QA dataset, and CAMeL-2 Naous and Xu (2025), an extractive QA dataset featuring culturally grounded entities. BLEnD Myung et al. (2024) is a hand-crafted benchmark designed to evaluate LLMs’ everyday knowledge across diverse cultures in English. It comprises 500 short-answer question-answer pairs for each country, where the answers vary depending on the country’s cultural or regional context. To reduce computational cost, we exclude North Korea and West Java, resulting in a final selection of 14 countries from the BLEnD dataset.
CAMeL-2 Naous and Xu (2025) is a bilingual benchmark originally constructed to evaluate LLMs’ entity extraction capabilities on Arabic and English entities. With CAMeL-2, an LLM is asked to extract an entity from a context collected from Arabic X/Twitter data according to the specified entity type in the input. We take the English partition and reduce the dataset to 14 countries to keep a similar country distribution to the BLEnD dataset. Dataset details, including domains and countries covered by the datasets, can be found in Appendix A.
3.2.2 Grouping of Cultures
We categorize 14 countries from each dataset along the resource dimension and the region dimension to study how overgeneralization manifests across these dimensions. For the resource dimension, we adopt the taxonomy proposed by Joshi et al. (2020), which classifies languages into six levels (0: very low-resource to 5: very high-resource). For our experiments, we simplify this into three groups: High (Level 5), Mid (Levels 3-4), and Low (Levels 0-2). We assign each country a language resource level based on its most widely spoken language provided by Wikipedia. For the region dimension, we group countries into six regions based on continents. We split Asia into three subregions, which leaves us six regional groups: North America, Europe, Africa, West Asia, South Asia, and East Asia. A complete list of countries within each group is provided in Appendix A.1.
3.2.3 Cultural MCQ with hard negatives
Khan et al. (2025) found that if MCQs lack the adversarial depth to probe genuine cultural understanding, models can exploit surface-level elimination strategies without truly understanding cultural distinctions. Thus, we propose a cultural MCQ with hard negatives to study how overgeneralization—driven by regional or resource dominance or similarity—affects the downstream task. Since BLEnD Myung et al. (2024) provides different answers from each culture with the same question, we create BLEnD-resource and BLEnD-region partition using culturally nuanced answers in BLEnD.
Figure 2: A cultural question about the popular family game in China from the BLEnD dataset Myung et al. (2024). For the given question about China, if an LLM answers ‘Yutnori’, a popular family game in South Korea, it is caused by the effect of cultural flattening between South Korea and China. On the other hand, if the answer is ‘Monopoly’, the LLM is generating an answer from a high resource culture.
We design two types of multiple choice question option lists that incorporate hard negative options: $C_{resource}$ and $C_{region}$, corresponding to BLEnD-resource and BLEnD-region, respectively. For $C_{resource}$, given a question $q$ targeting country $y$, we sample one country from each of the three resource levels excluding $y$. We obtain these three country’s respective gold answers when substituted into $q$ for $y$, resulting in three hard negative options: $c_{high}$, $c_{mid}$, and $c_{low}$. For $C_{region}$, we sample one country from the same geographical region as $y$ (excluding $y$) and extract its corresponding answer to construct a region-based hard negative option, $c_{flatten}$. Two additional options are randomly selected from countries in different regions. All options are shuffled to avoid positional bias. Figure 2 shows the example of $C_{resource}$ and $C_{region}$ for the question about China’s popular family game. Cultural MCQ with hard negative options allows us to examine when a model generates incorrect answers to cultural questions, whether a model’s cultural confusion arises from similarity in resource level or regional proximity.
3.3 Models
Application of MI methods requires a full-access to the model weights. Due to the requirement, we conduct our experiments with three recent open-sourced LLMs: Meta-Llama-3.1-8B-Instruct (Llama-3.1, Grattafiori et al. (2024)), aya-expanse-8b (aya-expanse, Dang et al. (2024)), and Qwen2.5-7B-Instruct (Qwen2.5, Qwen Team (2024)).
3.4 Patchscope
Patchscope Ghandeharioun et al. (2024) utilizes an LLM itself to generate natural language explanations of its internal representations. It consists of two forward passes, one using a source prompt and the other using an inspection prompt, with a patching operation between them. An inspection prompt is designed to guide an LLM as a probe to extract specific knowledge encoded in its internal representations, aligned with a predefined objective, such as next-token prediction, or attribute extraction. Utilizing an LLM itself with an inspection prompt as a probing mechanism addresses key limitations of prior methods Hernandez et al. (2024); Geva et al. (2022); Belrose et al. (2025), which often rely on predefined probing classes or suffer from limited interpretability due to sub-word tokenization.
In the context of cultural knowledge, these limitations are particularly pronounced. It is challenging to exhaustively define all relevant cultural knowledge. Additionally, a direct projection onto the vocabulary space via the unembedding matrix becomes difficult to interpret since cultural knowledge is frequently tokenized into multiple tokens Naous and Xu (2025). To address these challenges, we introduce a Patchscope-based method tailored for probing the cultural knowledge space. To our knowledge, this is the first work to apply interpretability techniques for investigating cultural knowledge in LLMs.
4 Probing Cultural Knowledge within Internal Layers
Probing the cultural knowledge processed by each layer for the given input provides insights into how cultural knowledge for one culture is overlapping with different cultures within the inner layers of an LLM. To translate internal representations of LLMs to natural language that reveals the cultural knowledge space, we propose Culturescope, building upon the existing interpretability method, Patchscope Ghandeharioun et al. (2024).
4.1 Culturescope
Culturescope consists of three stages: inference, scoping-in, and filtering. Our Culturescope allows us to move beyond what is observable from model responses alone, overcoming the limitation of extrinsic evaluation.
Step 1. Inference
An LLM first encodes a tokenized input $T_{i}$ of $i$-th instance and generates an output $O_{i}$, which is an LLM answer to an Open-Ended cultural QA. Then, we compute the representative hidden representations used to generate an answer, which are patched onto the inspection prompt during the scoping-in stage.
As Patchscope does not consider patching with multi-tokens Ghandeharioun et al. (2024), we adopt Bronzini et al. (2024)’s approach originally developed for fact-checking claims to condense an LLM’s cultural answer involving multiple tokens into a single hidden representation. Similarly, we compute the representative hidden representation $x_{*}^{l}$ of the $O_{i}$, which consists of multiple tokens, on the layer $l$. Specifically, we perform the weighted sum of hidden states as in Eq. 2 for the layer $l$. We set the weight $w_{p}$ of each token to one if it is a noun or a verb. Other token weights are set to zero. We use this $x_{*}^{l}$ in the next stage.
$x_{*}^{l}=\sum_{p=0}^{P}x_{p}^{l}*w_{p}$ (2)
Step 2. Scoping-in
At this stage, we further elicit the cultural knowledge encoded in $x_{*}^{l}$ to reveal the cultural knowledge space utilized for the $O_{i}$. To elicit a list of cultural knowledge $\mathcal{CK}_{i}=[ck_{i,1},...,ck_{i,j},$...] from the cultural knowledge space, we design an inspection prompt to elicit cultural knowledge by computing a forward pass. The inspection prompt ends with a placeholder token ‘x’, where we perform the patching as in the Patchscope Ghandeharioun et al. (2024). At the $l$-th layer, we replace the hidden representation at the placeholder token position with $x_{*}^{l}$. The inspection prompt can be found in Appendix B.
Step 3. Filtering
We empirically observe that an LLM tends to generate knowledge that is not culture-specific with our inspection prompt when the patched representation lacks the cultural knowledge. Since our method aims at eliciting any cultural knowledge available within inner representations, we devise a filtering method rather than identifying the most relevant model component.
To filter out the knowledge that is unrelated to cultural knowledge, we calculate the semantic similarity as an activation score between the input text $T$ and the generated cultural knowledge $ck_{i,j}$. We take a DeBERTA natural language inference (NLI) model to obtain the hidden representation. [^2] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 In Eq. 3, we compute the representative representation of the input text $T$ using the final hidden states from the NLI model. The activation score $s_{i,j}$ of $ck_{i,j}$ is calculated by their cosine similarity.
$g_{t}^{*}=\frac{1}{T}\sum_{t=0}^{T}g_{t}$ (3)
We keep $ck_{i,j}$ when its $s_{i,j}$ is higher than the threshold, which is set to 0.3.
4.2 Cultural Flattening Score
LLMs trained on imbalanced cultural resources are likely to represent less-documented cultures through the cultural knowledge of more dominant ones, potentially leading to overgeneralization. To quantify this phenomenon, we introduce a Cultural Flattening score (CF score) that measures the extent to which one country’s learned representation has been homogenized to resemble another’s.
A CF score is asymmetric and calculated for a pair of countries, target country $y_{t}$ and source country $y_{s}$. To calculate the CF score, we first compute a cultural knowledge signature for each country. A cultural knowledge signature is a collection of cultural knowledge decoded from Culturescope (§4.1). Let $\mathcal{CK}$ denote the set of cultural knowledge and $\mathcal{Y}$ the set of all countries. For a country $y\in\mathcal{Y}$, let $S_{y}(ck)$ represent the set of activation scores obtained by Eq. 3 for knowledge $ck\in\mathcal{CK}_{y}$ across all instances from country $y$. We define the unnormalized knowledge signature for country $y$ as:
$\tilde{\sigma}_{y}(ck)=\begin{cases}\bar{s}_{y,ck}\cdot\log(1+|S_{y}(ck)|)&\text{if }S_{y}(ck)\neq\emptyset\\ 0&\text{otherwise}\end{cases}$ (4)
where $\bar{s}_{y,ck}=\frac{1}{|S_{y}(ck)|}\sum_{s\in S_{y}(ck)}s$ is the average activation score for knowledge $ck$ in country $y$, and $|S_{y}(ck)|$ denotes the frequency of knowledge $ck$ appearing in country $y$’s data.
The normalized cultural knowledge signature is then computed as:
$\sigma_{y}(ck)=\frac{\tilde{\sigma}_{y}(ck)}{\sum_{ck^{\prime}\in\mathcal{CK}}\tilde{\sigma}_{y}(ck^{\prime})}$ (5)
We define the set of universal concepts as those appearing in every country: $\mathcal{U}=\{ck\in\mathcal{CK}:\forall y\in\mathcal{Y},S_{y}(ck)\neq\emptyset\}$. The CF score from target country $y_{t}$ to source country $y_{s}$ is:
$F(y_{t}\rightarrow y_{s})=\sum_{ck\in\mathcal{CK}\setminus\mathcal{U}}\sigma_{y_{t}}(ck)\cdot\mathbb{1}[S_{y_{s}}(ck)\neq\emptyset]$ (6)
where $\mathbb{1}[\cdot]$ is the indicator function. This formulation ensures that $F(y_{t}\rightarrow y_{s})\in[0,1]$, with higher values indicating that a larger fraction of $y_{t}$’s culturally distinctive knowledge is shared with $y_{s}$.
This asymmetric nature of scores directs cultural influence: a high score $F(y_{t}\rightarrow y_{s})$ indicates that country $y_{t}$’s representation substantially overlaps with country $y_{s}$’s, suggesting that the model may have learned to represent $y_{t}$ through the $y_{s}$’s cultural knowledge.
5 Tracing LLMs’ Internal Mechanisms for Cultural Knowledge Usage
Leveraging MCQs with hard negatives (§3.2.3), we propose to investigate how LLMs internalize Western-dominance bias and cultural flattening via the attention map Yuksekgonul et al. (2024). This setup facilitates an analysis of whether cultural biases are also reflected in the attention mechanism, as in extrinsic evaluation, tracing the internal mechanisms for the emergence of cultural biases.
5.1 Attention Contribution Score
To examine how LLMs internally process Western-dominance bias and cultural flattening, we analyze how attention patterns are directed toward each option $c$ in the input $T$ using the MCQs with hard negatives which contains option list simulating the bias (§3.2.3). Following Yuksekgonul et al. (2024)’s work that highlights the final input token as a meaningful anchor point for attention analysis, we track attention from this final input token $t_{s}$ to tokens that correspond to options in an option list $C$.
For a tokenized input text $T$ that contains a culture-specific question $q$ and a curated option list, we compute the attention contribution $a_{t_{c},t_{s}}$ from the final input token $t_{s}$ to the token $t_{c}$ corresponding to an option $c$, using the following procedure. The operation involves four projection matrices $W^{L}_{Q},W^{L}_{K},W^{L}_{V},W^{L}_{O}\in\mathbb{R}^{d\times d}$ that correspond to the ‘query’, ‘key’, ‘value’, and ‘output’ projections in the attention block of the layer. Each of these is split into multiple heads, where $W^{l,h}_{Q},W^{l,h}_{K},W^{l,h}_{V}\in\mathbb{R}^{d\times d_{h}}$ and $W^{l,h}_{O}\in\mathbb{R}^{d_{h}\times d}$ denote the matrices for each head $h$. $H$ is the total number of heads, $d_{h}$ is the dimensionality for each head. Embeddings are split into equal parts such that $d_{H}=\frac{d}{H}$. The attention weight matrix $A^{l,h}$, calculated as in Eq. 7, is comprised of the attention weight values computed by the attention head $h$ at layer $l$ over an input $T$ containing $S$ number of tokens, where Softmax is taken row-wise. The layer-wise attention contribution score, $a_{t_{c},t_{s}}^{l}$, is defined as in Eq. 8, where $A^{l,h}_{t_{c},t_{s}}$ is the attention weight of the token $t_{c}$ and $t_{s}$ from the matrix $A^{l,h}$.
$A^{l,h}=\mathrm{Softmax}(\frac{(X^{l-1}W^{l,h}_{Q})(X^{l-1}W^{l,h}_{K})^{T}}{\sqrt{d_{h}/H}})$ (7)
$a_{t_{c},t_{s}}^{l}=\sum_{h=1}^{H}A^{l,h}_{t_{c},t_{s}}(x^{l-1}_{t_{c}}W^{l,h}_{V})W^{l,h}_{O}$ (8)
The final attention contribution score, $a_{t_{c}}$, is computed by averaging the layer-wise attention contribution scores $a_{t_{c},t_{s}}^{l}$ across all layers in the LLM, where $t_{s}$ is the final input token.
6 Experimental Results
BLEnD
Llama-3.1 aya-expanse Qwen2.5
Baseline 0.4848 0.4683 0.4626
Cultural Prompting 0.4699 0.4638 0.4664
CANDLE 0.4007 0.2692 0.3473
Culturescope 0.5462 0.4928 0.5059
CAMeL-2
Llama-3.1 aya-expanse Qwen2.5
Baseline 0.7126 0.7019 0.6675
Cultural Prompting 0.6799 0.7282 0.7148
CANDLE 0.5640 0.6955 0.6473
Culturescope 0.6519 0.7169 0.6659
Table 1: We present the QA performances of three different LLMs with different inputs on BLEnD Myung et al. (2024) and CAMeL-2 Naous and Xu (2025). We highlight the best performing method in bold, and the second-best in italics.
6.1 Open-Ended QA Performances
To ensure its relevance to the input, we prepend the comma-separated cultural knowledge $CK$ to the input text $T$. The input example can be found in Appendix B. Given the augmented input $\mathcal{CK}_{i};T_{i}]$, an LLM is asked to generate an output $O$. To evaluate an LLM’s answer $O$, we perform exact-match evaluation, whether $O$ contains a gold answer $a$. We compare our method to three different input schemes, which are designed to enhance LLMs’ cultural understanding ability. Table 1 shows the accuracy from different input schemes.
Baseline is an input without an explicit instruction and additional cultural knowledge. Cultural Prompting Li et al. (2024); Cheng et al. (2023) is an input with an additional instruction (e.g. “The following question is about country yy”) designed to guide an LLM with an explicit country name. CANDLE Nguyen et al. (2023b) is a comprehensive cultural commonsense knowledge base consisting of triples (country, topic, assertion) spanning 196 countries. It provides cultural concepts derived from assertions, which we compare to the cultural knowledge decoded from Culturescope.
Since topics in CANDLE and those in the datasets are not aligned, we sample 20 concepts for each country. Due to the sampling, we perform the inference with three different random seeds and report the average accuracy. Experimental results show that inputs augmented with Culturescope perform better than those with CANDLE concepts. The accuracy with Culturescope is best on BLEnD across all models, on par on CAMeL-2 dataset with the baseline. This confirms that Culturescope reveals a valid cultural knowledge space, which is highly relevant to the given input.
Figure 3: We present the results from the CF score on BLEnD with Llama-3.1. A connection from $y_{t}$ to $y_{s}$ means that the fraction of $y_{t}$’s distinctive knowledge results from the cultural flattening with $y_{s}$.
6.2 Results with CF Score
Figure 3 illustrates the country pairs that exhibit cultural flattening, as measured by our CF score. To highlight only the most significant connections, we apply a threshold to exclude pairs with CF scores lower than the average of all country pairs. Notably, Assam and Indonesia are omitted from the visualization due to their relatively low scores. Overall, the results suggest that countries within the same regional groups tend to share a common cultural knowledge space with bidirectional connections. Additionally, Iran, the United States, and China appear prominently as source countries ($y_{s}$), indicating their broader influence on the connected target countries ($y_{t}$). Further results using CAMeL-2 and other models are available in Appendix C.
BLEnD-Resource BLEnD-Region
Acc % Biased % Others Refusal Acc % Biased % Others Refusal
Llama-3.1 avg. 0.43 0.19 0.18 0.02 0.43 0.20 0.18 0.02
mid 0.44 0.19 0.18 0.02 0.45 0.21 0.16 0.02
low 0.39 0.21 0.19 0.02 0.37 0.19 0.21 0.03
aya-expanse avg. 0.38 0.18 0.16 0.11 0.35 0.19 0.17 0.13
mid 0.40 0.17 0.16 0.11 0.37 0.20 0.15 0.13
low 0.31 0.20 0.19 0.12 0.30 0.19 0.19 0.13
Qwen2.5 avg. 0.44 0.20 0.16 0.05 0.44 0.20 0.16 0.05
mid 0.44 0.19 0.17 0.04 0.47 0.19 0.15 0.04
low 0.40 0.21 0.17 0.06 0.36 0.21 0.18 0.06
Table 2: Model outputs result from Llama-3.1, aya-expanse, and Qwen2.5 on the BLEnD-Resource and BLEnD-Region dataset (§3.2.3), evaluated using four metrics. Results are aggregated by question’s target country type: all averaged (avg.), mid-resource (mid), and low-resource (low).
6.3 Performances on MCQ with Hard Negatives
Table 2 shows the model’s final output results from Llama-3.1, aya-expanse, and Qwen2.5 on the BLEnD-Resource and BLEnD-Region (§3.2.3). We aggregate the results by question’s target country type: all averaged (avg.), mid-resource (mid), and low-resource (low). This reflects LLM’s output-level preference for hard negative options over other options, which serves to reveal how internal biases within the LLM can affect its final prediction. We present the percentage of instances where the model chooses a gold answer, denoted as Accuracy (Acc). The metric labeled as “% Biased” indicates the proportion of instances where the model chooses a hard negative option, which represents a targeted bias. “% Others” represents the proportion of instances where the model chooses one of the remaining random options. Since there are two random options present in the option list, we divide the proportion of choosing random options by two for a fair comparison. We also report the proportion of instances as ‘Refusal’ where LLMs avoid answering.
In most cases, LLMs prefer a hard negative option to other options when they are generating wrong answers with a higher % Biased than % Others. By breaking down the results by resource levels Joshi et al. (2020), we observe that the accuracy decreases for low-resource target questions compared to the average. We also find that in BLEnD-resource, the proportion for selecting a hard negative option from high-resource cultures increases, suggesting the presence of Western-dominance bias. However, this trend does not necessarily extend to BLEnD-region, where the selection rate of hard negatives does not show a similar increase. This indicates that low-resource cultures may be less susceptible to cultural flattening, likely due to the limited cultural knowledge available in the model.
(a) BLEnD-Resource
(b) BLEnD-Region
Figure 4: We present a heatmap visualization of attention contribution scores (z-score normalized) for the Llama-3.1 model on incorrect predictions. Groups on x-axis represent the group of the country which the gold answer is from. Groups on y-axis represent the group of the country which the chosen option is from. For example, in Figure 4 (a), ‘Low (x-axis)’-‘High (y-axis)’ pair with the score of 0.44 shows the averaged attention contribution score on option tokens from ‘High’ resource group when the gold answer is from ‘Low’ resource group. In BLEnD-Resource, we observe high attention contribution scores to tokens from high-resource cultures. BLEnD-Region also shows similar trends, higher scores to Europe and North America, regional group consisting of high-resource countries.
6.4 Attention Contribution Score Analysis
Figure 4 presents the average of attention contribution scores on option token positions, assigned by Llama-3.1 when the model makes incorrect predictions. We separate the analysis between correct predictions and incorrect predictions, as we are particularly interested in LLMs’ internal patterns when they are making biased predictions. Details for the aggregation method and results from aya-expanse and Qwen2.5 are shown in Appendix D.
In both BLEnD-Resource and BLEnD-Region, groups on x-axis represent groups of the question’s target country—resource level (High, Mid, Low) in BLEnD-Resource and region (South Asia [S-AS], East Asia [E-AS], West Asia [W-AS], Europe [EUR], America [AME], Africa [AFR]) in BLEnD-Region — while groups on y-axis indicate groups of the country which the chosen option is from. In Figure 4 (a), visualizing results of BLEnD-resource, we find that attention contributions from the last input token to incorrect high-resource country options are higher than incorrect mid- and low-resource countries options, especially for low-resource target questions. Figure 4 (b), visualizing results of BLEnD-region, shows similar trends to Figure 4 (a). Llama-3.1 allocates higher attention contributions to European and North American countries options compared to other region groups options. This trend is also shown in other two models as in Figure 7.
The analysis of attention contributions demonstrates that high-resource bias and Western-dominance bias is highly internalized within LLMs’ representations. This contradicts the evaluation on LLMs’ performance in Table 2, where LLMs are equally exhibiting both biases. This further aligns with findings from societal biases where they find intrinsic bias metrics and extrinsic bias metrics do not always correlate Goldfarb-Tarrant et al. (2021).
7 Discussion
7.1 Cultural Flattening within LLMs’ Inner Representations
With our Culturescope method, we can now probe the cultural knowledge encoded within the internal representations of LLMs. As shown in Figure 3, which visualizes the cultural flattening direction between cultures, we find unidirectional connections that have Iran and the United States as $y_{s}$. This unidirectional connection implies that models may have learned to represent less documented cultures, such as Ethiopia and Algeria, through those high-resource cultures. Our experiments using hard negative options align with previous works, which find that LLMs sometimes respond with answers aligned with culturally similar or geographically proximate regions Cao et al. (2023); Tao et al. (2024). We further attribute the models’ tendency to favor culturally adjacent answers to the unidirectional connections found by the proposed Culturescope. These findings underscore the need for methods that can disentangle culturally entangled representations, particularly among similar cultures, to enhance the accuracy and cultural appropriateness of LLM outputs.
7.2 LLMs’ Performance on Low-Resource Cultures
Previous studies Azime et al. (2025); Myung et al. (2024); Li et al. (2024) have shown that LLMs often struggle to utilize knowledge relevant to low-resource cultures. Our MCQ results with hard-negative options (Table 2; §6.3) are consistent with these findings, reinforcing the narrative that LLMs underperform on cultural tasks involving low-resource regions. Additionally, the relatively low cultural flattening (CF) scores for low-resource cultures (§6.2) suggest that these models are less prone to generate culturally flattened outputs due to the limited cultural knowledge encoded in their parameters. This is further supported by the reduced ratio of biased answers with region-type hard negatives for low-resource cultures, as shown in Table 2. These findings indicate that LLMs exhibit weaker cultural biases toward low-resource cultures, not because of improved fairness, but due to a lack of cultural representation. Consequently, improving performance on low-resource cultures may require a different approach — one that prioritizes knowledge acquisition over bias mitigation.
8 Conclusion
In this work, we investigate the complex and often biased ways LLMs process cultural knowledge. We introduce Culturescope, a method that leverages activation patching to probe the internal mechanisms of LLMs, allowing us to analyse the cultural knowledge encoded within their layers. In addition, we quantify the phenomenon of ‘cultural flattening’, where LLMs represent less-documented cultures through the concepts of more dominant or geographically close ones, thereby erasing cultural nuances. We conduct our research across three distinct models using culturally-grounded QA datasets, moving beyond isolated extrinsic evaluations to examine the interaction between cultural knowledge within the models’ parametric space. Our analysis reveals that models over-attend Western-centric tokens internally, indicating internalized Western-dominance bias. The results for low-resource cultures suggest that they may not be as susceptible to cultural biases as other cultures. Our findings suggest that future work should develop a tailored approach that considers the impact of bias and resource levels to improve LLMs’ cultural understanding.
Limitations
While our study provides new insights into cultural knowledge in LLMs, there exists limitations. Although a growing number of benchmarks aim to evaluate cultural knowledge in LLMs, few are suitable for our evaluation setup. To meaningfully compare the probed cultural knowledge across cultures, the datasets must maintain a consistent QA format across different cultures. In addition, to ensure that the evaluation captures how LLMs represent culture on a global scale, it is essential to include a geographically diverse set of cultures. However, we emphasize that our proposed method is model- and task-agnostic, and can be applied to any dataset that meets these requirements.
Due to computational constraints, we are reporting results with  8B models, unable to conduct experiments on larger-scale models. In addition, our analysis does not cover all countries globally. This limitation is not due to methodological oversight but rather the lack of publicly available datasets that include culturally grounded QA data for many regions. As such, our findings are necessarily constrained by the scope of existing resources. Finally, our MCQ with hard negatives (§3.2.3) involves a degree of random sampling. While this introduces some variability, we consider it a reasonable trade-off given the prohibitive cost of exhaustively evaluating all possible negative combinations. We mitigate this by ensuring consistency across runs and focusing on aggregate trends rather than individual instances.
Acknowledgements
[…]
Appendix A Dataset Details
Dataset Task Number of Questions Domain List of Countries
BLEnD
Myung et al. (2024)
Cultural
Commonsense QA
5726
Education, Food, Holidays/Celebration/Leisure,
Sport, Work life, Family
Africa: Algeria, Ethiopia, Northern Nigeria
Europe: Spain, United Kingdom, Greece
North America: United States, Mexico
East Asia: China, South Korea
South Asia: Indonesia, Assam
West Asia: Iran, Azerbaijan
CAMeL-2
Naous and Xu (2025)
Extractive QA 1862 Locations, Beverage, Food, Sports
Africa: Morocco, Algeria
Europe: Spain, United Kingdom, Greece
North America: United States, Mexico
East Asia: China, Japan
South Asia: Indonesia, India
West Asia: Iran, Syria, Egypt
Table 4: Details for the datasets
In our experiments, we utilize BLEnD Myung et al. (2024) and CAMeL-2 Naous and Xu (2025). We provide their brief data statistics and characteristics in Table 4.
A.1 Country Groups
As mentioned in Section 3.2, our work conducts experiments that focus on 14 countries classified in two dimensions. We compare three groups based on the level of language resource Joshi et al. (2020) and six groups based on the continental region. Table 3 shows the country entities that correspond to each group.
Appendix B Example of Prompts
In Figure 5, we present the prompt templates we use for each method to obtain Table 1.
Appendix C CF Score Results
In Figure 6, we show the CF score results with Llama-3.1, aya-expanse, Qwen2.5 on BLEnD and CAMeL-2. As mentioned in §6.2, we exclude the countries with CF scores lower than the average CF score across all countries.
(a) BLEnD, aya-expanse
(b) BLEnD, Qwen2.5
(c) CAMeL-2, aya-expanse
(d) CAMeL-2, Qwen2.5
(e) CAMeL-2, Llama-3.1
Figure 6: CF score results
Appendix D Attention Contributions
In cases where an option consists of multiple tokens, we follow the approach of Yuksekgonul et al. (2024), taking the maximum attention contribution score among the component tokens. When the attention scores are averaged by samples to examine its general patterns, simply averaging can be sensitive to extreme values or samples in which attention scores are globally biased toward higher or lower magnitudes. This variability can potentially reduce the generalizability of the results. To address this, we applied normalization (z-score normalization) per sample to the attention contribution scores, such that the scores within each sample have a mean of 0 and a standard deviation of 1.
(a) BLEnD-Resource - Qwen2.5-7B-Instruct
(b) BLEnD-Region - Qwen2.5-7B-Instruct
(c) BLEnD-Resource - Aya Expanse 8B
(d) BLEnD-Region - Aya Expanse 8B
Figure 7: Heatmap visualization of average attention contribution scores (z-score normalized) on incorrect predictions.